Find two numbers whose sum is 7 and whose difference is 23.
Answer by blwinbbbles(106)

Start by writing equations..We know that:
x+y=7
x-y=23 we can use the elimination process by adding the two equations together and solving
so:
2x = 30
x = 15
Now put it back in and solve:
15 + y = 7
y = -8
So your answer is (15,-8)